OpenStudy (angiepangie0726):

Solve 2x2 + 16x + 27 = 0 by completing the square.

OpenStudy (angiepangie0726):

@welshfella

OpenStudy (angiepangie0726):

I got x equals negative four plus or minus square root of start fraction five over two end fraction end square root

OpenStudy (4everaddicted2anime):

First divide everything by 2 to get the x^2 by itself

OpenStudy (angiepangie0726):

is my answer right?

OpenStudy (welshfella):

first divide through by 2 x^2 + 8x + 27/2= 0 (x + 4)^2 - 16 + 13.5 = 0 (x+ 4)^2 = 2.5 x + 4 = +/- sqrt2.5

OpenStudy (welshfella):

x = -4 +/- sqrt2.5

OpenStudy (welshfella):

yes you are right ( 5/2 = 2.5)